# Nonlinear Schrödinger-Poisson systems in dimension two: the zero mass case

### v1

@@ Abstract.
We provide an existence result for a Schrödinger-Poisson system in gradient form, set in the whole plane, in the case of zero mass. Since the setting is limiting for the Sobolev embedding, we admit nonlinearities with subcritical or critical growth in the sense of Trudinger-Moser. In particular, the absence of the mass term requires a nonstandard functional framework, based on homogeneous Sobolev spaces. These features, combined with the logarithmic behaviour of the kernel of the Poisson equation, make the analysis delicate, since standard variational tools cannot be applied. The system is solved by considering the corresponding logarithmic Choquard equation. We prove the existence of a mountain pass-type solution via a careful analysis on specific Cerami sequences, whose boundedness is achieved by exploiting an appropriate functional, obtained by evaluating the energy functional on particular paths.
@@ 1. Introduction
Substituting it in the first equation of (1.2), one may rewrite the system as a Choquard equation, that is an integro-differential equation of Schrödinger type with a convolutive right-hand side:
@@
Besides the evident advantage of the reduction to a single equation, since the system (1.2) is of gradient type, if $N\geq 3$ the equation (1.3) is variational in the Sobolev space $H^{1}(\mathbb{R}^{N})$ thanks to the Hardy-Littlewood-Sobolev inequality (Proposition 2.5 below), provided suitable polynomial growth conditions on $f$ are fulfilled. In this respect, there are a huge number of works about Choquard-type equations, especially from the last decades, and we refer to the seminal works [35, 36, 16, 15] and the references therein. The planar case $N=2$ is more delicate because of the interplay between the logarithmic behaviour of the Riesz kernel, and the exponential maximal growth of the nonlinearities, due to the Pohožeav-Trudinger-Moser inequality in the full space proved by Ruf [42], see also [8]. The first attempt of considering this case is to be referred to Stubbe [45] and later on formalised by Cingolani and Weth [19, 20]: they set the problem in a constraint space which takes into account in the seminorm a contribution of the logarithmic kernel. This analysis, which is peculiar for the case of a linear coupling in the system, namely $f(u)=u$ in (1.2), was then extended for the general case of a nonlinearity with critical exponential growth in [14]. Taking indeed into account the behaviour at 00 of the nonlinearity, and by means of a log-weighted version of the Pohožaev-Trudinger inequality, a proper functional setting was found, in which the functional associated to (1.3) turns out to be well-defined. We also point out that the sharp version of such inequality has been recently obtained in [46]. The approach in [14] was then generalised for Choquard equations with weights in [9] and for quasilinear Schrödinger-Poisson systems in [7]. A different approach was recently proposed in [33]: here, instead, the underlying functional space remains $H^{1}(\mathbb{R}^{2})$, while the logarithmic kernel is uniformly approximated by polynomial kernels. For further developments of this method we refer to [11, 13, 12], also in quasilinear fractional contexts. All the above works deal with Choquard equations of the form (1.3), where $V$ is a nontrivial potential. The special case of an identically zero potential, the so-called “zero-mass case”, emerges in some physical context, e.g. in the nonabelian gauge theory of particle physics, such as the study of the Yang-Mills equation, see [27]. The main difference with the “mass-case” lies in the natural framework in which such problems are studied, namely the homogeneous Sobolev spaces $D^{1,2}_{0}(\mathbb{R}^{N})$ defined as the completion of $C^{\infty}_{c}(\mathbb{R}^{N})$ with respect of the $L^{2}$-norm of the gradient. In the higher-dimensional case $N\geq 3$, in light of the critical Sobolev embedding into $L^{2^{*}}(\mathbb{R}^{N})$ with $2^{*}=\frac{2N}{N-2}$, such setting is appropriate for both Schrödinger and Choquard equations of the kind
@@
as shown e.g. in [5, 3, 1, 2]. However, if $N=2$, not only any embedding into Lebesgue spaces is out of reach, but also one cannot distinguish in $D^{1,2}_{0}(\mathbb{R}^{2})$ between functions which just differ by constants. In this framework only few results are available for Schrödinger equations and Choquard equations of the kind (1.4) with $K_{2}=\frac{1}{2\pi}\ln\frac{1}{|\cdot|}$. On the one hand, one may try to tweak the operator, so that the corresponding natural space recovers good embedding properties in Lebesgue and Orlicz spaces: this is the strategy used in [21, 10] for the Schrödinger case, and in [40, 41, 22] for the Choquard case with both polynomial and logarithmic kernels. On the other hand, for equations with zero mass driven by the pure Laplacian, the results available in the literature [47, 18, 17] cover just the linear case $f(u)=u$ in (1.4) (up to adding local nonlinearities) since the approach of [45, 19] is pursued. However, as remarked in [17],
@@
Indeed, differently from the Schrödinger equations with zero mass, the presence of the logarithmic kernel combined with $f(u)=u$ allows to recover $H^{1}(\mathbb{R}^{2})$ as suitable functional framework, thanks to a careful splitting in positive and negative part of the logarithm (see (3.2) below).
Main goal of this paper is to extend the existence results in the zero-mass case of [47, 18, 17] in the direction of [14], that is aiming at considering the general case of a nonlinear function $f$, and covering both the cases of subcritical and critical growth in the sense of Ruf’s inequality. To this aim, several difficulties need to be faced: first, the unusual functional setting, which does not appear in the above cited works because of the linear behaviour of $f$; then the analysis on Cerami sequences, which arise from the mountain pass geometry of the functional associated to (1.5), is largely affected by the possibly exponential growth of the nonlinear terms and, in particular in the critical case, it is very delicate; eventually also the final proof “à la Lions” of the existence theorem is pretty non-standard and rather technical. Finally, we derive from our results for the Choquard equation the corresponding for the Schrödinger-Poisson system (1.1) in a suitable functional setting in the spirit of [7, 40]. We stress that this step is often neglected in the literature by just considering it as “natural”: here it finds a rigorous justification.
Before stating our main results, let us specify the growth conditions we are assuming on the nonlinearity. Throughout the paper we suppose that $f\in C^{1}(\mathbb{R})$, $f(s)>0$ as $s>0$, while $f(s)=0$ for $s\leq 0$; moreover it satisfies:
@@
    for any $\alpha>0$, $\lim_{s\to+\infty}f(s)/{\rm e}^{\alpha s^{2}}=0$ and for some $C>0$, $f(s)\geq Cs^{p-1}$ as $s\to+\infty\,$;
@@
$D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2}):=D^{1,2}_{0}(\mathbb{R}^{2})\cap L^{p}(\omega\,{\rm d}x)\,,$
where the weight function $\omega(x):=\ln(b+|x|)$ with $b>1$. Note that this space, which corresponds to the intersection space $H^{1}L^{p}_{\omega}(\mathbb{R}^{2})$ detected in [14] for the study of the same Choquard equation with positive mass, gathers all important features of our problem: the absence of mass since we are dealing with the homogeneous Sobolev space $D^{1,2}(\mathbb{R}^{2})$, the nonlinear behaviour of $f$, and the logarithmic kernel in the weight. We will see that this space enjoys good embedding properties, in particular exponential nonlinearities are allowed. We are therefore lead to the following definition:
@@
    $f^{\prime}(s)\asymp s^{p-2}$ as $s\to 0$ and there exists $\alpha^{\prime}>0$ and $s^{\prime}>0$ such that $f^{\prime}(s)\lesssim{\rm e}^{\alpha^{\prime}s^{2}}$ for $s>s^{\prime}$.
We postpone to Section 2.3 a detailed list of consequences of our assumptions. Here we just emphasise that $(f_{2})$ implies an Ambrosetti-Rabinowitz condition and the monotonicity of $f$; $(f_{3})$-$(f_{4})$ will be used in the analysis of the boundedness of Cerami sequences when dealing with critical nonlinearities, the latter being related to the deFiguereido-Miyagaki-Ruf condition in [23] and used to prove a fine upper bound for the mountain pass level in Section 4; $(f_{5})$ is a mild condition about the growth at $\infty$ of the nonlinearity, which well agrees with both $(f_{sc})$-$(f_{c})$, and will be exploited in the conclusive compactness argument in Section 5. Our main result reads as follows.
@@
Once we have found a weak solution of the logarithmic Choquard equation (1.5), we can go back to the original Schrödinger-Poisson system. First, we need a precise meaning of solution for (1.1).
The weighted Lebesgue space $L_{s}(\mathbb{R}^{2})$, $s>0$, is defined as
@@
Under the conditions of Theorem 1.2, the Schrödinger-Poisson system (1.1) possesses a solution $(u,\Phi)\in D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})\times L_{s}(\mathbb{R}^{2})$ for all $s>0$ such that $u$ is positive and $\Phi=\Phi_{u}:=\ln\frac{1}{|\cdot|}\ast F(u)$.
@@
    it is sufficient to prove the existence of a nonnegative nontrivial solution of (1.5) in order to retrieve its positivity by the strong maximum principle for semilinear equations, see e.g. [38, Theorem 11.1].
@@
In Section 2 we describe the functional framework in which it is convenient to set our problem, discuss our assumptions, and collect some useful results. The variational framework is then described in Section 3, where we show the mountain pass geometry for the energy functional, while the existence of special Cerami sequences, and their boundedness is detailed in Section 4; we stress that these arguments turn out to be a delicate matter. After some careful mountain pass estimates, the proof of the existence for the log-Choquard equation (1.5) is given in Section 5. Finally, in Section 6 we derive from it the existence result for the Schrödinger-Poisson system (1.1).
Notation.
For $R>0$ and $x_{0}\in\mathbb{R}^{N}$ we denote by $B_{R}(x_{0})$ the ball of radius $R$ and center $x_{0}$. Given a set $\Omega\subset\mathbb{R}^{N}$, we denote $\Omega^{c}:=\mathbb{R}^{N}\setminus\Omega$, and its characteristic function by $\chi_{\Omega}$. The space of the infinitely differentiable functions which are compactly supported is denoted by $C^{\infty}_{c}(\mathbb{R}^{N})$, $\mathcal{M}(\mathbb{R}^{N})$ stands for the space of measurable functions in $\mathbb{R}^{N}$, while ${\mathcal{S}}$ is the Schwartz space of rapidly decreasing functions and ${\mathcal{S}}^{\prime}$ the dual space of tempered distributions. For $p\in[1,+\infty]$ the Lebesgue space of $p$-integrable functions is denoted by $L^{p}(\mathbb{R}^{N})$ with norm $\|\cdot\|_{p}$. For $q>1$ we define its conjugate Hölder exponent as $q^{\prime}:=\frac{q}{q-1}$. The symbol $\lesssim$ indicates that an inequality holds up to a multiplicative constant depending only on structural constants, while $f\asymp g$ means that $c_{1}f\leq g\leq c_{2}f$ for some $c_{1},c_{2}>0$. Finally, $\text{o}_{n}(1)$ denotes a vanishing real sequence as $n\to+\infty$. Hereafter, the letter $C$ will be used to denote positive constants which are independent of relevant quantities and whose value may change from line to line.
2. Functional space and preliminary results
2.1. The space $\boldsymbol{D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})}$
@@
$D^{1,2}(\mathbb{R}^{2}):=\left\{u\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{2})\,|\,\nabla u\in L^{2}(\mathbb{R}^{2})^{2}\right\}$
with seminorm $\|\nabla\cdot\|_{2}$. Note that, by the unboundedness of the domain, this seminorm cannot control the $L^{2}$-norm of the elements of $D^{1,2}(\mathbb{R}^{2})$, and therefore $D^{1,2}(\mathbb{R}^{2})\supsetneq H^{1}(\mathbb{R}^{2})$. To retrieve a normed space, one need to introduce the relation $u\sim v$Leftrightarrow $v=u+c$ with $c\in\mathbb{R}$, and define the quotient space $\dot{D}^{1,2}(\mathbb{R}^{2}):=\left\{[u]\,|\,u\in D^{1,2}\right\}$, which turns out to be a Hilbert space with norm $\|\nabla\cdot\|_{2}$ (see [26, Lemma II.6.2]). On the other hand, one may also introduce the space
@@
By [26, Theorem II.7.5] the two spaces are isomorphic: $D^{1,2}_{0}(\mathbb{R}^{2})=\dot{D}^{1,2}(\mathbb{R}^{2})$. In order to find a suitable variational framework for the system (1.1), for $p>2$ and $b>1$ we define
@@
Note that, by the choice of $b>1$, in $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$ it is possible to control the $p-$norm by the seminorm $\|u\|_{\ast,p}$. Indeed,
(2.1) $\|u\|^{p}_{p}=\int_{\mathbb{R}^{2}}|u(x)|^{p}\,{\rm d}x\leq(\ln b)^{-1}\int_{\mathbb{R}^{2}}\ln(b+|x|)|u(x)|^{p}\,{\rm d}x=(\ln b)^{-1}\|u\|^{p}_{*,p}\,.$
@@
and so, since $C^{\infty}_{c}(\mathbb{R}^{2})$ is dense both in $D^{1,2}_{0}(\mathbb{R}^{2})$ and $L^{p}(\mathbb{R}^{2})$, we have the characterisation
$D^{1,2}(\mathbb{R}^{2})\cap L^{p}(\mathbb{R}^{2}\omega\,{\rm d}x)$ $=D^{1,2}(\mathbb{R}^{2})\cap L^{p}(\mathbb{R}^{2})\cap L^{p}(\mathbb{R}^{2},\omega\,{\rm d}x)$
@@
Furthermore, by [30, Theorem 1.11] $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$ is a reflexive Banach space, whose dual is given by
(2.2) $D^{-1,2}L^{p}_{\omega}(\mathbb{R}^{2}):=(D^{1,2}(\mathbb{R}^{2})\cap L^{p}_{\omega}(\mathbb{R}^{2}))^{\prime}=D^{-1,2}(\mathbb{R}^{2})\big{|}_{D^{-1,2}L^{p}_{\omega}}+L^{p^{\prime}}(\mathbb{R}^{2},\omega\,{\rm d}x)\big{|}_{D^{-1,2}L^{p}_{\omega}},$
since $\left(L^{p}_{\omega}(\mathbb{R}^{2})\right)^{\prime}=L^{p^{\prime}}(\mathbb{R}^{2},\omega\,{\rm d}x)$, see [44, Theorem 14.9], and [26, Theorem II.8.1] for the representation of the space $D^{-1,2}(\mathbb{R}^{2})$. Let us state now important embedding properties of our space.
Lemma 2.1.
@@
The embedding $X\hookrightarrow L^{p}(\mathbb{R}^{2})$ is a consequence of (2.1). Let now $q>p$, then by Proposition 2.6 below (applied with $j=0,m=1,r=2,N=2,q=p$) we have
@@
Let us now prove the compactness of these embeddings by relying on the Riesz criterion, see [39, Theorem XIII.66], which needs the continuity of the translation in the Lebesgue norm and a uniform decay at infinity of the elements in $X$. To this aim, let $S\subset X$ be a bounded subset, then it is also bounded in $L^{q}(\mathbb{R}^{2})$, $q\geq p$. Let $R>0$, $u\in S$ and $q\geq p$. Then, by Hölder inequality,
$\int_{\{|x|\geq R\}}|u|^{q}\,{\rm d}x\leq\|u\|_{(q-1)p^{\prime}}^{q-1}\int_{\{|x|\geq R\}}|u|^{p}\,{\rm d}x\leq C\int_{\{|x|\geq R\}}|u|^{p}\,{\rm d}x,$
@@
since $u\in S$ bounded in $X$. Hence, for $q\geq p$, for any $\varepsilon>0$ one can choose $R>0$ large enough such that
(2.3) $\int_{B_{R}(0)^{c}}|u|^{q}\,{\rm d}x\leq\varepsilon^{q}.$
Let us now prove the continuity of the translation in $L^{q}(\mathbb{R}^{2})$. Since $X\subset D^{1,2}_{0}(\mathbb{R}^{2})$, by density we can work within $C^{\infty}_{c}(\mathbb{R}^{2})$. Let $u\in C^{\infty}_{c}(\mathbb{R}^{2})$ and $h\in\mathbb{R}^{2}$. Following [6, Proposition 9.3] and defining $\tau_{h}u:=u(\cdot+h)$, by Jensen inequality we have

### v2

@@ Abstract.
We provide an existence result for a Schrödinger-Poisson system in gradient form, set in the whole plane, in the case of zero mass. Since the setting is limiting for the Sobolev embedding, we admit nonlinearities with subcritical or critical growth in the sense of Trudinger-Moser. In particular, the absence of the mass term requires a nonstandard functional framework, based on homogeneous Sobolev spaces. These features, combined with the logarithmic behaviour of the kernel of the Poisson equation, make the analysis delicate, since standard variational tools cannot be applied. The system is solved by considering the corresponding logarithmic Choquard equation. The existence of a mountain pass-type solution is established by means of a careful analysis of appropriate Cerami sequences, whose boundedness is ensured through a nonstandard variational method, suggested by the subtle nature of the functional geometry involved. As a key tool in our estimates, we also introduce a logarithmic weighted Trudinger–Moser inequality, along with a related Cao-type inequality, both of which hold in our functional setting and are, we believe, of independent interest.
@@ 1. Introduction
Substituting it into the first equation of (1.2), one may rewrite the system as a Choquard equation, that is an integro-differential equation of Schrödinger type with a convolutive right-hand side:
@@
Besides the evident advantage of the reduction to a single equation, since the system (1.2) is of gradient type, if $N\geq 3$ the equation (1.3) is variational in the Sobolev space $H^{1}(\mathbb{R}^{N})$ thanks to the Hardy-Littlewood-Sobolev inequality (Proposition 2.9 below), provided suitable polynomial growth conditions on $f$ are fulfilled. In this respect, there are a huge number of works about Choquard-type equations, especially from the last decades, and we refer to the seminal works [36, 37, 16, 15] and the references therein. The planar case $N=2$ is more delicate because of the interplay between the logarithmic behaviour of the Riesz kernel, and the exponential maximal growth of the nonlinearities, due to the Pohožaev-Trudinger-Moser inequality in the full space proved by Ruf [43], see also [8]. The first attempt of considering this case is to be referred to Stubbe [46], later on formalised by Cingolani and Weth [19, 20]: they set the problem in a constrained space which takes into account in the seminorm a contribution of the logarithmic kernel. This analysis, which is peculiar for the case of a linear coupling in the system, namely $f(u)=u$ in (1.2), was then extended for the general case of a nonlinearity with critical exponential growth in [14]. Taking indeed into account the behaviour at 00 of the nonlinearity, and by means of a log-weighted version of the Pohožaev-Trudinger inequality, a proper functional setting was found, in which the functional associated to (1.3) turns out to be well-defined. We also point out that the sharp version of such inequality has been recently obtained in [47]. The approach in [14] was then generalised for Choquard equations with weights in [9] and for quasilinear Schrödinger-Poisson systems in [7]. A different approach was recently proposed in [34]: here, instead, the underlying functional space remains $H^{1}(\mathbb{R}^{2})$, while the logarithmic kernel is uniformly approximated by polynomial kernels. For further developments of this method we refer to [11, 12, 13], also in quasilinear fractional contexts. All the above works deal with Choquard equations of the form (1.3), where $V$ is a nontrivial potential. The special case of an identically zero potential, the so-called “zero-mass case”, emerges in some physical context, e.g. in the nonabelian gauge theory of particle physics, such as the study of the Yang-Mills equation, see [27]. The main difference with the “mass-case” lies in the natural framework in which such problems are studied, namely the homogeneous Sobolev spaces $D^{1,2}_{0}(\mathbb{R}^{N})$ defined as the completion of $C^{\infty}_{c}(\mathbb{R}^{N})$ with respect of the $L^{2}$-norm of the gradient. In the higher-dimensional case $N\geq 3$, in light of the critical Sobolev embedding into $L^{2^{*}}(\mathbb{R}^{N})$ with $2^{*}=\frac{2N}{N-2}$, such setting is appropriate for both Schrödinger and Choquard equations of the kind
@@
as shown e.g. in [5, 3, 1, 2]. However, if $N=2$, not only any embedding into Lebesgue spaces is out of reach, but also one cannot distinguish in $D^{1,2}_{0}(\mathbb{R}^{2})$ between functions which just differ by constants. In this framework only few results are available for Schrödinger equations and Choquard equations of the kind (1.4) with $K_{2}=\frac{1}{2\pi}\ln\frac{1}{|\cdot|}$. On the one hand, one may try to adjust the operator so that the corresponding natural space recovers good embedding properties in Lebesgue and Orlicz spaces: this is the strategy used in [21, 10] for the Schrödinger case, and in [42, 41, 22] for the Choquard case with both polynomial and logarithmic kernels. On the other hand, for equations with zero mass driven by the pure Laplacian, the results available in the literature [48, 18, 17] cover just the linear case $f(u)=u$ in (1.4) (up to adding local nonlinearities) since the approach of [46, 19] is pursued. However, as remarked in [17],
@@
Indeed, differently from the Schrödinger equations with zero mass, the presence of the logarithmic kernel combined with $f(u)=u$ allows to recover $H^{1}(\mathbb{R}^{2})$ as suitable functional framework, thanks to a careful splitting in positive and negative part of the logarithm (see (3.2) below). Main goal of this paper is to extend the existence results in the zero-mass case of [48, 18, 17] in the direction of [14], that is to consider the general case of a nonlinear function $f$, and cover both the cases of subcritical and critical growth in the sense of the Trudinger-Moser inequality. To this aim, several difficulties need to be faced: first, the unusual functional setting, which does not appear in the above cited works because of the linear behaviour of $f$; then the analysis on Cerami sequences, which arise from the mountain pass geometry of the functional associated to (1.5), is largely affected by the possibly exponential growth of the nonlinear terms and, in particular in the critical case, it is very delicate; eventually the final proof “à la Lions” of the existence theorem is also quite non-standard and rather technical. Finally, we derive from our results for the Choquard equation the corresponding for the Schrödinger-Poisson system (1.1) in a suitable functional setting in the spirit of [7, 42]. We stress that this step is often neglected in the literature by just considering it as “natural”: here it finds a rigorous justification. Before stating our main results, let us specify the growth conditions we are assuming on the nonlinearity. Throughout the paper, we suppose that $f\in C^{1}(\mathbb{R})$, $f(s)>0$ as $s>0$, while $f(s)=0$ for $s\leq 0$; moreover, it satisfies:
@@
    for any $\alpha>0$, $\lim_{s\to+\infty}f(s)/{\rm e}^{\alpha s^{2}}=0$ and for some $C>0$, $f(s)\geq Cs^{p-1}$ as $s\to+\infty\,$, with $p>2$;
@@
$D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2}):=D^{1,2}_{0}(\mathbb{R}^{2})\cap L^{p}(\mathbb{R}^{2},\omega\,{\rm d}x)\,,$
with weight function $\omega(x):=\ln(b+|x|)$ and $b>1$. Note that this space, which corresponds to the intersection space $H^{1}L^{p}_{\omega}(\mathbb{R}^{2})$ detected in [14] for the study of the same Choquard equation with positive mass, gathers all important features of our problem: the absence of mass since we are dealing with the homogeneous Sobolev space $D^{1,2}(\mathbb{R}^{2})$, the nonlinear behaviour of $f$, and the logarithmic kernel in the weight. We will see that this space enjoys good embedding properties, in particular exponential nonlinearities are allowed. We are therefore led to the following definition:
@@
    $f^{\prime}(s)\asymp s^{p-2}$ as $s\to 0$ with $p>2$, and $f^{\prime}$ satisfies ($f_{c}$) with the same $\alpha_{0}$ as $f$.
We postpone to Section 2.4 a detailed list of consequences of our assumptions. Here we just emphasise that $(f_{2})$ implies an Ambrosetti-Rabinowitz condition and the monotonicity of $f$; $(f_{3})$-$(f_{4})$ will be used in the analysis of the boundedness of Cerami sequences when dealing with critical nonlinearities, the latter being related to the de Figueiredo-Miyagaki-Ruf condition in [23] and used to prove a fine upper bound for the mountain pass level in Section 4; $(f_{5})$ is a mild condition about the growth at $\infty$ of the nonlinearity, which well agrees with both $(f_{sc})$-$(f_{c})$, and will be exploited in the conclusive compactness argument in Section 5. Our main result reads as follows.
@@
Once we have found a weak solution of the logarithmic Choquard equation (1.5), we can go back to the original Schrödinger-Poisson system. First, we need a precise meaning of solution for (1.1). The weighted Lebesgue space $L_{s}(\mathbb{R}^{2})$, $s>0$, is defined as
@@
Under the conditions of Theorem 1.2, the Schrödinger-Poisson system (1.1) possesses a solution $(u,\Phi)\in D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})\times L_{s}(\mathbb{R}^{2})$ for all $s>0$, such that $u$ is positive and $\Phi=\Phi_{u}:=\ln\frac{1}{|\cdot|}\ast F(u)$.
@@
    it is sufficient to prove the existence of a nonnegative nontrivial solution of (1.5) in order to retrieve its positivity by the strong maximum principle for semilinear equations, see e.g. [39, Theorem 11.1].
@@
In Section 2 we describe the functional framework in which it is convenient to set our problem, and prove Trudinger-Moser- and Cao-type inequalities in weighted spaces, which will be fundamental tools in our arguments; moreover, we discuss our assumptions and collect some useful results. The variational framework is then described in Section 3, where we show the mountain pass geometry for the energy functional, while the existence of special Cerami sequences, and their boundedness is detailed in Section 4; we stress that these arguments turn out to be a delicate matter. After some careful mountain pass estimates, the proof of the existence for the log-Choquard equation (1.5) is given in Section 5. Finally, in Section 6 we derive from it the existence result for the Schrödinger-Poisson system (1.1).
Notation
For $R>0$ and $x_{0}\in\mathbb{R}^{N}$ we denote by $B_{R}(x_{0})$ the ball of radius $R$ and centre $x_{0}$, and we omit the centre when $x_{0}=0$. Given $\Omega\subset\mathbb{R}^{N}$, we denote $\Omega^{c}:=\mathbb{R}^{N}\setminus\Omega$, and its characteristic function by $\chi_{\Omega}$. The space of infinitely differentiable functions which are compactly supported is denoted by $C^{\infty}_{c}(\mathbb{R}^{N})$, $\mathcal{M}(\mathbb{R}^{N})$ stands for the space of measurable functions in $\mathbb{R}^{N}$, while ${\mathcal{S}}$ is the Schwartz space of rapidly decreasing functions and ${\mathcal{S}}^{\prime}$ the dual space of tempered distributions. For $p\in[1,+\infty]$ the Lebesgue space of $p$-integrable functions is denoted by $L^{p}(\mathbb{R}^{N})$ with norm $\|\cdot\|_{p}$. For $q>1$ we define its conjugate Hölder exponent as $q^{\prime}:=\frac{q}{q-1}$. The symbol $\lesssim$ indicates that an inequality holds up to a multiplicative constant depending only on structural constants, while $f\asymp g$ that $c_{1}f\leq g\leq c_{2}f$ for some $c_{1},c_{2}>0$. Finally, $\text{o}_{n}(1)$ denotes a vanishing real sequence as $n\to+\infty$. Hereafter, the letter $C$ will be used to denote positive constants which are independent of relevant quantities and whose value may change from line to line.
2. Functional space, functional inequalities, and preliminaries
2.1. The space $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$
@@
$D^{1,2}(\mathbb{R}^{2}):=\left\{u\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{2})\,|\,\nabla u\in[L^{2}(\mathbb{R}^{2})]^{2}\right\}$
with seminorm $\|\nabla\cdot\|_{2}$. Note that, by the unboundedness of the domain, this seminorm cannot control the $L^{2}$-norm of the elements of $D^{1,2}(\mathbb{R}^{2})$, and therefore $D^{1,2}(\mathbb{R}^{2})\supsetneq H^{1}(\mathbb{R}^{2})$. To retrieve a normed space, one needs to introduce the relation $u\sim v$Leftrightarrow $v=u+c$ with $c\in\mathbb{R}$, and define the quotient space $\dot{D}^{1,2}(\mathbb{R}^{2}):=\left\{[u]_{\sim}\,|\,u\in D^{1,2}\right\}$, which turns out to be a Hilbert space with norm $\|\nabla\cdot\|_{2}$ (see [26, Lemma II.6.2]). On the other hand, one may also introduce the space
@@
By [26, Theorem II.7.5] the spaces $D^{1,2}_{0}(\mathbb{R}^{2})$ and $\dot{D}^{1,2}(\mathbb{R}^{2})$ are isomorphic. In order to find a suitable variational framework for the system (1.1), for $p>2$ and $b>1$ we define
@@
Note that, by the choice of $b>1$, in $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$ it is possible to control the $L^{p}-$norm by the seminorm $\|\cdot\|_{\ast,p}$. Indeed,
(2.1) $\|u\|^{p}_{p}\leq(\ln b)^{-1}\int_{\mathbb{R}^{2}}\ln(b+|x|)|u(x)|^{p}\,{\rm d}x=(\ln b)^{-1}\|u\|^{p}_{*,p}\,.$
@@
and so, since $C^{\infty}_{c}(\mathbb{R}^{2})$ is dense in both $D^{1,2}_{0}(\mathbb{R}^{2})$ and $L^{p}(\mathbb{R}^{2})$, we have the characterisation
$D^{1,2}(\mathbb{R}^{2})\cap L^{p}(\mathbb{R}^{2},\omega\,{\rm d}x)$ $=D^{1,2}(\mathbb{R}^{2})\cap L^{p}(\mathbb{R}^{2})\cap L^{p}(\mathbb{R}^{2},\omega\,{\rm d}x)$
@@
Furthermore, by [31, Theorem 1.11] $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$ is a reflexive Banach space, whose dual is given by
$D^{-1,2}L^{p}_{\omega}(\mathbb{R}^{2}):$ $=(D^{1,2}(\mathbb{R}^{2})\cap L^{p}_{\omega}(\mathbb{R}^{2}))^{\prime}$
$=D^{-1,2}(\mathbb{R}^{2})\big{|}_{D^{-1,2}L^{p}_{\omega}}+L^{p^{\prime}}(\mathbb{R}^{2},\omega\,{\rm d}x)\big{|}_{D^{-1,2}L^{p}_{\omega}},$
since $\left(L^{p}_{\omega}(\mathbb{R}^{2})\right)^{\prime}=L^{p^{\prime}}(\mathbb{R}^{2},\omega\,{\rm d}x)$, see [45, Theorem 14.9], and also [26, Theorem II.8.1] for the representation of the space $D^{-1,2}(\mathbb{R}^{2})$. Let us now state important embedding properties of our space.
Proposition 2.1.
@@
The embedding $X\hookrightarrow L^{p}(\mathbb{R}^{2})$ is a consequence of (2.1). Let now $q>p$, then by the Gagliardo-Nirenberg inequality (see Proposition 2.10 below, applied with $j=0,m=1,r=2,N=2,q=p$) we have
@@
Let us now prove the compactness of these embeddings by relying on the Riesz criterion, see [40, Theorem XIII.66], which requires the continuity of the translation in the $L^{q}-$norm and a uniform decay at infinity of the elements in $X$. To this aim, let $S\subset X$ be a bounded subset, which is then also bounded in $L^{q}(\mathbb{R}^{2})$ for $q\geq p$. Let also $R>0$ and $u\in S$. Then, by the Hölder inequality,
$\int_{\{|x|\geq R\}}|u|^{q}\,{\rm d}x\leq\|u\|_{(q-1)p^{\prime}}^{q-1}\left(\int_{\{|x|\geq R\}}|u|^{p}\,{\rm d}x\right)^{\frac{1}{p}}\leq C\left(\int_{\{|x|\geq R\}}|u|^{p}\,{\rm d}x\right)^{\frac{1}{p}},$
@@
since $u\in S$ bounded in $X$. Hence, for any $\varepsilon>0$ one can choose $R>0$ large enough such that
(2.2) $\int_{B_{R}^{c}}|u|^{q}\,{\rm d}x\leq\varepsilon^{q}.$
Let us now prove the continuity of the translation in $L^{q}(\mathbb{R}^{2})$. Since $X\subset D^{1,2}_{0}(\mathbb{R}^{2})$, by density we can work within $C^{\infty}_{c}(\mathbb{R}^{2})$. Let $u\in C^{\infty}_{c}(\mathbb{R}^{2})$ and $h\in\mathbb{R}^{2}$. Following [6, Proposition 9.3] and defining $\tau_{h}u:=u(\cdot+h)$, by the Jensen inequality we have